200 people attend a school concert, 3/4 of the audience were female. What percentage were male?
Found 2 solutions by richwmiller, amomoses:
Answer by richwmiller(17219)

Are we counting transgenders, bisexuals, and crossdressers as distinct genders?
If not then 25% of the audience is male.

Answer by amomoses(4)

3/4*100%=75%
girls%=75%
so boys%=100-75=25%